# WHAM: Reconstructing World-grounded Humans with Accurate 3D Motion

### v1

@@ 3.3 Training
To synthesize 2D keypoints from AMASS, we create virtual cameras onto which we project 3D keypoints derived from the ground truth mesh. Unlike MotionBERT [61] and ProxyCap [60], which use static cameras for keypoint projection, we employ dynamic cameras that incorporate both rotational and translational motion. This choice is based on two main reasons. First, it accounts for the inherent differences between human motion captured in static and dynamic camera setups. Second, it enables the learning of a camera-agnostic motion representation, from which the trajectory decoder can reconstruct the global trajectory. We also augment the 2D data with noise and masking. For details of the synthetic generation process see Sup. Mat.
Fine-tuning on Video Datasets. Starting with the pre-trained motion encoder and decoders, we fine-tune WHAM on four video datasets: 3DPW [49], Human3.6M [11], MPI-INF-3DHP [33], and InstaVariety [15]. For the human mesh recovery task, we supervise WHAM on ground-truth SMPL parameters from AMASS and 3DPW, 3D keypoints from Human3.6M and MPII3D, and 2D keypoints from InstaVariety. For the global trajectory estimation task, we use AMASS, Human3.6M, and MPII3D. Additionally, during training we experiment with adding BEDLAM [1](which we call WHAM-B), a large synthetic dataset with realistic video and ground truth SMPL parameters.
The fine-tuning has two objectives: 1) exposing the network to real 2D keypoints, instead of training it solely on synthetic data, and 2) training the feature integrator network to aggregate motion and image features. To achieve these goals, we jointly train the entire network on the video datasets while setting a smaller learning rate on the pre-trained modules (see Fig. 3). Consistent with prior work [17, 30, 6, 51, 43], we employ a pre-trained and fixed-weight image encoder [21] to extract image features. However, to leverage recent network architectures and training strategies, we also experiment with different types of encoders [25, 1, 7] in the following section.
@@ 4 Experiments
Datasets. We evaluate WHAM on three in-the-wild benchmarks: 3DPW [49], RICH [10], and EMDB [16]. Following previous work [14, 21, 17, 6, 30, 25, 1], we perform the evaluation in camera coordinates. The estimated global trajectory is evaluated on a subset of EMDB (EMDB 2) for which they provide ground truth global motion with dynamic cameras (used for evaluation). We also test on new sequences captured using an iPhone with ground-truth camera angular rotation from the gyroscope. See Sup. Mat. for more details of the datasets and iPhone results.
[…]
@@
per-frame SPIN [21]   59.2 96.9 112.8 31.4   69.7 122.9 144.2 35.2   87.1 140.3 174.9 41.3
$*{}^{*}$ [18]   46.5 74.5 88.6   60.7 109.2 123.5   72.2 113.9 133.2
$*{}^{*}$ [25]   43.0 69.0 81.2 22.5   56.6 102.6 115.0 22.4   68.1 103.3 128.0 24.5
$*{}^{*}$ [23]   41.8 71.6 82.3   56.4 96.8 110.4   65.6 103.0 122.2
HMR2.0 [7]   44.4 69.8 82.2 18.1   48.1 96.0 110.9 18.8   60.6 98.0 120.3 19.8
$*{}^{*}$ [1]   43.0 66.9 78.5 31.0   50.2 84.4 95.6 29.3   60.6 98.0 111.6 36.1
temporal PACE [20]     49.3 8.8
$*{}^{*}$ [6]   52.7 86.5 101.4 6.0   65.6 119.1 137.7 5.0   79.6 127.6 147.9 5.3
$*{}^{*}$ [17]   51.9 82.9 98.4 18.5   68.4 120.5 140.2 21.8   81.4 125.9 146.8 26.6
$*{}^{*}$ [51]   52.1 84.3 99.0 6.5   67.1 118.2 136.7 5.8   81.3 123.1 138.4 6.2
$*{}^{*}$ [43]   50.6 80.7 96.4 6.0   65.6 114.3 132.7 5.2   78.8 119.7 138.4 5.4
GLAMR [57]   51.1 8.0   79.9 107.7   73.5 113.6 133.4 32.9
$*{}^{*}$ [45]   50.9 79.1 95.4 28.6     70.9 109.9 127.4 25.5
SLAHMR [54]   55.9   52.5 9.4   69.5 93.5 110.7 7.1
WHAM ($*{}^{*}$   41.7 65.7 78.7 6.6   53.1 91.4 105.6 5.3   58.9 90.3 106.1 5.7
WHAM ($*{}^{*}$   40.9 64.5 77.7 6.8   50.7 88.0 100.8 5.7   58.4 89.5 107.7 6.5
WHAM ($*{}^{*}$   37.8 60.8 72.5 6.8   46.2 84.1 95.5 5.5   53.4 87.3 102.9 5.7
WHAM-B ($*{}^{*}$   37.2 59.4 71.0 6.9   44.7 82.6 93.2 5.6   48.8 80.7 93.7 5.9
Table 1: Quantitative comparison of state-of-the-art models on the 3DPW [49], RICH [10], and EMDB [16] datasets. Ordering of per-frame and temporal methods is done separately by descending MPJPE on EMDB. For testing on EMDB, we follow the protocol of EMDB 1 [16]. Parenthesis denotes the number of body joints used to compute MPJPE and PA-MPJPE, and $*{}^{*}$ denotes models trained with the 3DPW training set. Bold numbers denote the most accurate method in each column. Accel is in $m/s^{2}$, all other errors are in $mm$.
@@
Per-frame accuracy. In Table 1, we present a comprehensive comparison of WHAM and the existing state-of-the-art per-frame and video-based methods across three benchmark datasets [49, 10, 16]. WHAM (Res), WHAM (HR), and WHAM (ViT) correspond to different architectures for the pre-trained image encoders, derived from SPIN (ResNet-50) [21], CLIFF (HRNet-W48) [25, 1], and HMR2.0 (ViT-H/16) [7], respectively. Not surprisingly, WHAM (HR) is more accurate than WHAM (Res), while the transformer-based version, WHAM (ViT), is the most accurate. The backbone matters, with WHAM (ViT) outperforming all previous methods on all per-frame metrics (MPJPE, PA-MPJPE, and PVE) on all benchmarks. Because none of the methods are exposed to data from RICH or EMDB during training, results on these datasets are indicative of each method’s ability to generalize. Even with the simplest ResNet backbone, WHAM (Res) outperforms every method with the exception of BEDLAM-CLIFF on RICH. Unlike BEDLAM-CLIFF, WHAM (Res) is not trained on the BEDLAM dataset. Note that training on BEDLAM consistently improves accuracy in prior work (see [3, 1]), and we find the same here when we add BEDLAM to the training data (WHAM-B). SLAHMR results for 3DPW and RICH are from [54], while EMDB results are computed with their released code.
@@
Inter-frame smoothness. We also evaluate the inter-frame smoothness using the acceleration error. Compared with state-of-the-art per-frame methods [7, 25, 1, 23], WHAM has significantly lower acceleration error. This indicates that WHAM reconstructs smooth and more plausible 3D human motion across frames while not sacrificing high per-frame accuracy. Conversely, when compared to recent temporal methods [43, 51, 6], WHAM exhibits marginally higher acceleration error. However, we observe that these video-based methods tend to over-smooth the human motion, resulting in lower accuracy on per-frame metrics.
To provide intuition for these numbers, we qualitatively compare WHAM with TCMR [6] and GLoT [43] in Fig. 4. While producing smooth results, TCMR and GLoT fail to capture the bending of the left knee when the subject is ascending the stairs, while WHAM more accurately reconstructs the 3D human pose. For more qualitative results please see the Supplemental Video.
@@
[…]
DPVO (+ HMR2.0) [47, 7]   49.6 2320.9 662.9 17.5 44.4 112.8
GLAMR [57]   56.0 756.1 286.2 16.7 74.9 18.0
TRACE [45]   58.0 2244.9 544.1 18.9 72.7 370.7
SLAHMR [54]   61.5 807.4 336.9 13.8 67.9 19.7
WHAM (w/ DPVO [47])   41.9 446.6 169.0 8.8 40.4 14.8
WHAM (w/ DROID [46])   41.9 439.2 166.1 8.4 36.0 14.7
WHAM (w/ GT gyro)   41.9 436.4 165.9 7.1 26.3 14.8
@@
4.2 3D Global Trajectory Recovery
As shown in Table 2, WHAM outperforms the existing methods on all metrics. Specifically, combining WHAM with DPVO is more accurate than the global trajectory estimation of DPVO combined with HMR2.0, illustrating that our method actively refines the global trajectory instead of performing a simple integration. DROID-SLAM gives slightly better results than DPVO. Furthermore, WHAM significantly outperforms TRACE on egocentric root and foot velocities. We further demonstrate this in Figs. 5 and WHAM: Reconstructing World-grounded Humans with Accurate 3D Motion, where WHAM captures more consistent and plausible human motion in the global coordinate system than TRACE and SLAHMR for videos captured by dynamic cameras. As depicted in Fig. 6, WHAM outperforms GLAMR, TRACE, and SLAHMR in capturing the pattern of human motion in the global coordinate system. See the Supplemental Video for more examples.
@@ 4.3 Ablation Study
Methods   PA-MPJPE PVE RTE ROE ERVE EFVE
w/o $F_{I}$   44.1 84.9 9.4 41.6 15.1 20.7
w/o lifting   108.9 187.5 18.5 98.3 29.4 36.8
w/o traj. ref.   41.9 74.7 9.0 40.4 14.4 20.9
w/o $\omega$   40.7 74.5 18.3 77.1 14.4 20.8
WHAM (Ours)   41.9 74.7 8.8 40.4 14.8 19.8
@@
To provide further insight into our approach, we conduct ablation studies to analyze the contribution of each component to the performance. As shown in Table 3, our entire system (WHAM) outperforms the different variants of WHAM that ablate a single component. To be specific, first, we observe that adding feature integration improves both motion and global trajectory estimation accuracy when compared with an ablated version without feature integration (w/o $F_{I}$). Similarly, the removal of the pre-training on the 2D-to-3D lifting task using AMASS [32] (w/o lifting) shows significant performance degradation. In addition, we observe that WHAM without the trajectory refinement (w/o traj. ref.) gives higher root translation and foot velocity error, indicating that our refinement approach contributes to improving the global trajectory estimation and helps reduce foot sliding. Last, we experiment with WHAM to decode trajectory solely based on the motion context without using the estimated camera angular velocity (w/o $\omega$). Although this version shows robust performance on predicting 3D human pose, it suffers from the entanglement of camera and human motion, resulting in significantly high global trajectory errors (RTE and ROE).
@@ 5 Conclusion
WHAM is a new method to recover accurate 3D human motion in global coordinates from a moving camera more efficiently and accurately than the state-of-the-art approaches. Our approach leverages the AMASS dataset to train a network to recursively lift 2D sequences of keypoints to sequences of 3D SMPL parameters. But keypoints alone lack valuable information about the body and its movement. Consequently, we integrate image context information over time and learn to combine it with the motion context to better estimate human body shape and pose. Additionally, our method takes an estimate of the camera angular velocity, which can either be computed from a SLAM method or from the camera’s gyro when available. Finally, we combine all this information with an estimate of foot contact to recover the 3D human motion in global coordinates from a monocular video sequence. WHAM significantly outperforms the existing state-of-the-art methods (both image-based and video-based) on challenging in-the-wild benchmarks in both 3D motion and the world-coordinate trajectory estimation accuracy. Because of its speed and accuracy, WHAM provides a foundation for in-the-wild motion capture applications.
Limitations: WHAM learns about human motion from AMASS, limiting generalization to motions that are out of distribution. For example, since AMASS does not contain people riding bicycles or skateboards, WHAM does not capture global motion in these cases. Our contact estimation only applies to the feet and should be extended to include other body parts that may be in contact with the scene. Since WHAM relies on an estimate of the camera’s angular velocity, errors in this estimate can accumulate over time, leading to drift in the global trajectory. While we employ random masking as part of our data synthesis process, our generating approach mainly assumes the scenario where the full body is within the field of view. This can be addressed with additional augmentation during training (cf. [18, 26]). See the Supplemental Video. for example failure cases.
@@
Future directions: WHAM opens up many directions for future work. For example, while we use SLAM to estimate the camera’s angular velocity, SLAM could also provide camera intrinsics and extrinsics as well as information about the 3D scene that could be used to enforce consistency between the scene and the human. While WHAM is an online method, designed for real-time applications, it could also initialize an optimization-based post-processing akin to bundle adjustment, which would optimize the camera, scene, and human motion together. Furthermore, a real-time and phone-based implementation of WHAM should be feasible.
6 Acknowledgement
@@
We would like to sincerely appreciate Hongwei Yi and Silvia Zuffi for the discussion and proofreading. Part of this work was done when the first author was an intern at the Max Planck Institute for Intelligence System. This work was partially supported through an NSF CAREER Award (CBET 2145473) and a Chan-Zuckerberg Essential Open Source Software for Science Award.
@@
Disclosure. MJB has received research gift funds from Adobe, Intel, Nvidia, Meta/Facebook, and Amazon. MJB has financial interests in Amazon, Datagen Technologies, and Meshcapade GmbH. While MJB is a consultant for Meshcapade, his research in this project was performed solely at, and funded solely by, the Max Planck Society.
Appendix A Supplementary Materials
As promised in the main paper, this supplemental document provides details of our synthetic data generation, the datasets we use, our network training, and the run-time cost. Additionally, please refer to our Supplemental Video for results that illustrate our method and recent SOTA methods applied to video sequences.
A.1 Synthetic data generation

### v2

@@ 3.3 Training
To synthesize 2D keypoints from AMASS, we create virtual cameras onto which we project 3D keypoints derived from the ground truth mesh. Unlike MotionBERT [62] and ProxyCap [61], which use static cameras for keypoint projection, we employ dynamic cameras that incorporate both rotational and translational motion. This choice has two main motivations. First, it accounts for the inherent differences between human motions captured in static and dynamic camera setups. Second, it enables the learning of a camera-agnostic motion representation, from which the trajectory decoder can reconstruct the global trajectory. We also augment the 2D data with noise and masking. For details of the synthetic generation process see Sup. Mat.
Fine-tuning on Video Datasets. Starting with the pretrained network, we fine-tune WHAM on four video datasets: 3DPW [49], Human3.6M [11], MPI-INF-3DHP [33], and InstaVariety [15]. For the human mesh recovery task, we supervise WHAM on ground-truth SMPL parameters from AMASS and 3DPW, 3D keypoints from Human3.6M and MPI-INF-3DHP, and 2D keypoints from InstaVariety. For the global trajectory estimation task, we use AMASS, Human3.6M, and MPI-INF-3DHP. Additionally, during training we experiment with adding BEDLAM [1], a large synthetic dataset with realistic video and ground truth SMPL parameters.
The fine-tuning has two objectives: 1) exposing the network to real 2D keypoints, instead of training it solely on synthetic data, and 2) training the feature integrator network to aggregate motion and image features. To achieve these goals, we jointly train the entire network on the video datasets while setting a smaller learning rate on the pretrained modules (see Fig. 3). Consistent with prior work [17, 30, 6, 52, 43], we employ a pretrained and fixed-weight image encoder [21] to extract image features. However, to leverage recent network architectures and training strategies, we also experiment with different types of encoders [25, 1, 7] in the following section.
Implementation Details. For the pretraining stage, we train WHAM on AMASS for 80 epochs with the learning rate of $5\times 10^{-4}$. Then we finetune WHAM on 3DPW, MPI-INF-3DHP, Human3.6M, and InstaVariety for 30 epochs. We use a learning rate of $1\times 10^{-4}$ for the feature integrator and $1\times 10^{-5}$ for the pretrained components during finetuning. During training, we use the Adam optimizer and a batch size of 64 and split sequences into 81-frame chunks.
@@ 4 Experiments
Datasets. We evaluate WHAM on three in-the-wild benchmarks: 3DPW [49], RICH [10], and EMDB [16]. Following previous work [14, 21, 17, 6, 30, 25, 1], we perform the evaluation in camera coordinates. The estimated global trajectory is evaluated on a subset of EMDB (EMDB 2) for which they provide ground truth global motion with dynamic cameras (used for evaluation). We also test on new sequences captured using an iPhone with the gyroscope. See Sup. Mat. for more details on the datasets and iPhone results.
Evaluation metrics. To evaluate the accuracy of 3D human pose and shape estimation, we compute Mean Per Joint Position Error (MPJPE), Procrustes-aligned MPJPE (PA-MPJPE), and Per Vertex Error (PVE) measured in millimeters $(mm)$. We compute Acceleration error (Accel, in $m/s^{2}$) [^1] Previous work follows [15] in reporting Accel in $mm/\mathrm{frame}^{2}$. To remove the dependency on frame rate, we convert all previous results to $m/s^{2}$. to measure the inter-frame smoothness of the reconstructed motion. We also evaluate the motion reconstruction and trajectory estimation accuracy in the world-frame. Following previous work [55, 20], we split sequences into smaller segments of 100 frames and align each output segment with the ground-truth data using the first two frames (W-MPJPE100) or the entire segment (WA-MPJPE100) in $mm$. These previous metrics give an unrealistic picture of 3D performance as they do not measure drift over long sequences. Therefore, we also evaluate the error over the entire trajectory after the rigid alignment and measure Root Translation Error (RTE in %) normalized by the actual displacement of the person. We also assess the jitter of the motion in the world coordinate system in $10m/s^{3}$ and foot sliding during the contact (FS in $mm$).
@@
per-frame SPIN [21]  59.2 96.9 112.8 31.4    69.7 122.9 144.2 35.2    87.1 140.7 166.1 41.3
PARE [18]  46.5 74.5 88.6 –    60.7 109.2 123.5 –    72.2 113.9 133.2
CLIFF [25]  43.0 69.0 81.2 22.5    56.6 102.6 115.0 22.4    68.3 103.5 123.7 24.5
HybrIK [23]  41.8 71.6 82.3 –    56.4 96.8 110.4 –    65.6 103.0 122.2
HMR2.0 [7]  44.4 69.8 82.2 18.1    48.1 96.0 110.9 18.8    60.7 98.3 120.8 19.9
ReFit [51]  40.5 65.3 75.1 18.5    47.9 80.7 92.9 17.1    58.6 88.0 104.5 20.7
temporal TCMR [6]  52.7 86.5 101.4 6.0    65.6 119.1 137.7 5.0    79.8 127.7 150.2 5.3
VIBE [17]  51.9 82.9 98.4 18.5    68.4 120.5 140.2 21.8    81.6 126.1 149.9 26.5
MPS-Net [52]  52.1 84.3 99.0 6.5    67.1 118.2 136.7 5.8    81.4 123.3 143.9 6.2
GLoT [43]  50.6 80.7 96.4 6.0  65.6 114.3 132.7 5.2    79.1 119.9 140.8 5.4
GLAMR [58]  51.1 8.0    79.9 107.7    73.8 113.8 134.9 33.0
TRACE [45]  50.9 79.1 95.4 28.6    –    71.5 110.0 129.6 25.5
SLAHMR [55]  55.9 –    52.5 9.4    69.7 93.7 111.3 7.1
PACE [20]  –    49.3 8.8
WHAM (Res)  40.2 62.7 75.1 6.3    51.8 89.5 103.2 5.0  57.8 84.0 99.7 5.2
WHAM (HR)  39.0 62.6 74.8 6.4    49.1 84.6 96.4 5.2    57.1 85.7 103.2 5.6
WHAM (ViT)  35.9 57.8 68.7 6.6    44.3 80.0 91.2 5.3    50.4 79.7 94.4 5.3
Table 1: Quantitative comparison of state-of-the-art models on the 3DPW [49], RICH [10], and EMDB [16] datasets. Ordering of per-frame and temporal methods is done separately by descending MPJPE on EMDB (except for PACE). For testing on EMDB, we follow the protocol of EMDB 1 [16]. Parenthesis denotes the number of body joints used to compute MPJPE and PA-MPJPE, and denotes models trained with the 3DPW training set. Bold numbers denote the most accurate method in each column. Accel is in $m/s^{2}$, all other errors are in $mm$.
  3DPW (14)
Models Dataset    PA-MPJPE MPJPE PVE Accel
CLIFF [25, 1] R    43.6 68.8 82.1 19.2
R+B    43.0 66.9 78.5 31.0
ReFit [51] R    40.5 65.3 75.1 18.5
R+B    38.2 57.6 67.6 21.4
WHAM (ViT) R    35.9 57.8 68.7 6.6
R+B    35.7 56.9 67.4 6.7
Table 2: Dataset ablation experiments on 3DPW [49]. R denotes the use of real datasets and B denotes BEDLAM.
@@
Per-frame accuracy. Table 1 presents a comprehensive comparison of WHAM and the existing state-of-the-art per-frame and video-based methods across three benchmark datasets [49, 10, 16]. Because none of the methods are exposed to data from RICH or EMDB during training, results on these datasets are indicative of each method’s ability to generalize. WHAM (Res), WHAM (HR), and WHAM (ViT) correspond to different architectures for the pretrained image encoders, derived from SPIN (ResNet-50) [21], CLIFF (HRNet-W48) [25, 1], and HMR2.0 (ViT-H/16) [7], respectively. Not surprisingly, WHAM (HR) is more accurate than WHAM (Res), while the transformer-based version, WHAM (ViT), is the most accurate. The backbone matters, with WHAM (ViT) outperforming all previous methods on all per-frame metrics (MPJPE, PA-MPJPE, and PVE) on all benchmarks. Even with the simplest ResNet backbone, WHAM (Res) outperforms every method except for ReFit on RICH.
Training on BEDLAM consistently improves accuracy in prior work (see [3, 1, 51]), and we find the same here as shown in Table 2. However, compared to ReFit, WHAM exhibits relatively smaller performance improvement by adding BEDLAM. One possible reason for this is that, unlike ReFit, we did not finetune the image encoder during training to limit training time.
@@
Inter-frame smoothness. We also evaluate the inter-frame smoothness using the acceleration error. Compared with state-of-the-art per-frame methods [7, 25, 1, 23], WHAM has significantly lower acceleration error. This indicates that WHAM reconstructs smooth and more plausible 3D human motion across frames while not sacrificing high per-frame accuracy. On the other hand, when compared to recent temporal methods [43, 52, 6], WHAM exhibits comparable or slightly higher acceleration error. However, we observe that these video-based methods tend to over-smooth the human motion, resulting in lower accuracy on per-frame metrics.
To provide intuition for these numbers, we qualitatively compare WHAM with TCMR [6] and GLoT [43] in Fig. 4. While producing smooth results, TCMR and GLoT fail to capture the bending of the left knee when the subject is ascending the stairs, while WHAM more accurately reconstructs the 3D human pose.
@@
Models    WA-MPJPE100 W–MPJPE100 RTE Jitter FS
DPVO (+ HMR2.0) [47, 7]  647.8 2231.4 15.8 537.3 107.6
GLAMR [58]  280.8 726.6 11.4 46.3 20.7
TRACE [45]  529.0 1702.3 17.7 2987.6 370.7
SLAHMR [55]  326.9 776.1 10.2 31.3 14.5
WHAM (w/DPVO [47])  135.6 354.8 6.0 22.5 4.4
WHAM (w/DROID [46])  133.3 343.9 4.6 21.5 4.4
WHAM (w/ GT gyro)  131.1 335.3 4.1 21.0 4.4
@@
4.2 3D Global Trajectory Recovery
As shown in Table 3, WHAM outperforms the existing methods on all metrics. Specifically, combining WHAM with DPVO is more accurate than the global trajectory estimation of DPVO combined with HMR2.0, illustrating that our method actively refines the global trajectory instead of performing a simple integration. DROID-SLAM gives slightly better results than DPVO. Furthermore, WHAM significantly outperforms the regression-based method, TRACE, on jitter and foot sliding metrics. We further demonstrate this in Figs. 5 and 1, where WHAM captures more consistent and plausible human motion in the global coordinate system than TRACE and SLAHMR for videos captured by dynamic cameras. As depicted in Fig. 6, WHAM outperforms GLAMR, TRACE, and SLAHMR in capturing the pattern of human motion in the global coordinate system.
@@ 4.3 Ablation Study
To provide further insight into our approach, we conduct ablation studies to analyze the contribution of each component to the performance. As shown in Table 4, our entire system (WHAM) outperforms the different variants of WHAM that ablate a single component. To be specific, we first observe that adding feature integration improves both motion and global trajectory estimation accuracy when compared with an ablated version without feature integration (w/o $F_{I}$). Similarly, the removal of the pretraining on the 2D-to-3D lifting task using AMASS [32] (w/o lifting) shows significant performance degradation. WHAM also outperforms the ablation of the Neural Initialization (w/o NI), particularly in the MPJPE metric. In addition, we experiment with WHAM to decode trajectory solely based on the motion context without using the estimated camera angular velocity (w/o $\omega$). Although this version shows similar performance in predicting 3D human pose, it suffers from the entanglement of camera and human motion, resulting in significantly high overall trajectory error (RTE). Last, we observe that WHAM without the trajectory refinement (w/o traj. ref.) gives larger global trajectory and foot sliding errors in return for less jitter, indicating that our approach contributes to the global trajectory accuracy and helps reduce foot sliding. Core details are presented here; see Sup. Mat. for more details and information on run-time cost.
@@
Models    PA–MPJPE MPJPE WA–MPJPE100 W–MPJPE100 RTE Jitter FS
w/o $F_{I}$  44.2 69.0 147.6 377.9 6.3 23.1 5.5
w/o lifting    60.3 83.0 238.0 693.0 11.5 24.5 5.0
w/o NI  40.4 66.3 142.7 368.1 6.8 22.3 4.6
w/o $\omega$  39.1 62.0 156.5 422.0 10.1 22.1 5.0
w/o traj. ref.    38.2 59.3 154.7 407.5 6.3 18.8 6.5
WHAM (Ours)    38.2 59.3 135.6 354.8 6.0 22.5 4.4
@@ 5 Conclusion
WHAM is a new method to recover accurate 3D human motion in global coordinates from a moving camera more efficiently and accurately than the state-of-the-art approaches. Our approach leverages the AMASS dataset to train a network to recursively lift 2D sequences of keypoints to sequences of 3D SMPL parameters. But keypoints alone lack valuable information about the body and its movement. Consequently, we integrate image context information over time and learn to combine it with the motion context to better estimate human body shape and pose. Additionally, our method takes an estimate of the camera angular velocity, which can either be computed from a SLAM method or from the camera’s gyro when available. Finally, we combine all this information with an estimate of foot contact to recover the 3D human motion in global coordinates from a monocular video sequence. WHAM significantly outperforms the existing state-of-the-art methods (both image-based and video-based) on challenging in-the-wild benchmarks in both 3D HPS and the world-coordinate trajectory estimation accuracy. Because of its speed and accuracy, WHAM provides a foundation for in-the-wild motion capture applications.
Limitations and future directions: WHAM learns about human motion from AMASS, limiting generalization to motions that are out of distribution. While we employ random masking as part of our data synthesis process, our generating approach mainly assumes the scenario where the full body is within the field of view. See Sup. Mat. for more details.
@@
WHAM opens up many directions for future work. For example, while we use SLAM to estimate the camera’s angular velocity, SLAM could also provide camera intrinsics and extrinsics as well as information about the 3D scene that could be used to enforce consistency between the scene and the human. While WHAM is an online method, designed for real-time applications, it could also initialize an optimization-based post-processing akin to bundle adjustment, which would optimize the camera, scene, and human motion together. Furthermore, a real-time and phone-based implementation of WHAM should be feasible.
@@
Acknowledgements. Part of this work was done when the first author was an intern at the Max Planck Institute for Intelligence System. This work was partially supported through an NSF CAREER Award (CBET 2145473) and a Chan-Zuckerberg Essential Open Source Software for Science Award.
@@
CoI Disclosure. MJB has received research gift funds from Adobe, Intel, Nvidia, Meta/Facebook, and Amazon. MJB has financial interests in Amazon and Meshcapade GmbH. While MJB is a co-founder and Chief Scientist at Meshcapade, his research in this project was performed solely at, and funded solely by, the Max Planck Society.
Appendix A Synthetic data generation